OpenStudy (anonymous):

What is the distance between points (8, –4) and (8, –11) on a coordinate plane? units

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

medal and fan

OpenStudy (thehollywoodstud):

well do you know the equation to solve this?

OpenStudy (thehollywoodstud):

you have to make a triangle out of the two points like so

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ 8}}\quad ,&{\color{blue}{ -4}})\quad % (c,d) &({\color{red}{ 8}}\quad ,&{\color{blue}{ -11}}) \end{array}\qquad % distance value d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (thehollywoodstud):

8-8 is what ?

OpenStudy (anonymous):

0

OpenStudy (thehollywoodstud):

so 0+-11--4?

OpenStudy (solomonzelman):

| -4-(-11) |

OpenStudy (thehollywoodstud):

well you cant have two negatives so thats a positive

OpenStudy (thehollywoodstud):

so -11+4

OpenStudy (thehollywoodstud):

is what?

OpenStudy (thehollywoodstud):

which is 7

OpenStudy (jdoe0001):

anyhow, the fellow kinda dashed

OpenStudy (thehollywoodstud):

so then 7+0=7 do the d=sqrt(7)

OpenStudy (thehollywoodstud):

my bad then you have to do\[7^2=49 and the 4^2= 16 then add the two =65\]

OpenStudy (thehollywoodstud):

which is 8.06

OpenStudy (thehollywoodstud):

is that and answer choice ?@njrevine